# A chaotic lattice field theory in two dimensions

### v2

@@ VII Reciprocal lattice
Assembling the reciprocal primitive vectors $\{\tilde{\mathbf{a}}_{j}\}$ into columns of the $d\!\times\!d]$ reciprocal primitive cell matrix $\tilde{\mathbb{A}}=\left[\tilde{\mathbf{a}}_{1},\tilde{\mathbf{a}}_{2},\cdots,\tilde{\mathbf{a}}_{d}\right],$ the reciprocity condition Eq. (80) takes form
@@
VII.1 Reciprocal primitive cell in one and two dimensions
Figure 7: (Color online) (a) The intersection points $z$ of the light grey lines form the integer square lattice Eq. (6). The primitive vectors $\mathbf{a}_{1}=(3,0)$ and $\mathbf{a}_{2}=(1,2)$ form the primitive cell $\mathbb{A}={[3\!\times\!2]_{1}}$ (see Eq. (76) and Fig. 2 (a)), whose translations tile the Bravais lattice $\mathcal{L}_{\mathbb{A}}$ (red points). (b) The intersection points $k$ of the light grey lines form the reciprocal square lattice. Translations of reciprocal primitive vectors $\tilde{\mathbf{a}}_{1}$ and $\tilde{\mathbf{a}}_{2}$ (see Eq. (80), Eq. (89)) generate the reciprocal lattice $\mathcal{L}_{\tilde{\mathbb{A}}}$ (red points). (Shaded) The reciprocal primitive cell $\tilde{\mathbb{A}}$. A wave vector outside this region is equivalent to a wave vector within it by a reciprocal lattice translation. Note that the number of lattice sites within the reciprocal primitive cell $\tilde{\mathbb{A}}$ equals the number of sites within the spatiotemporal primitive cell $\mathbb{A}$.
@@ VII.1 Reciprocal primitive cell in one and two dimensions
with eigenvalues $\{{e}^{\mathrm{i}{k}}\}$ the n{n}-th roots of unity, indexed by integers $m$,
@@
$(a)$    Refer to caption
$(b)$    Refer to caption
$(c)$    Refer to caption
Figure 8: (Color online)    A one-dimensional temporal lattice period-5 periodic state $\mathsf{\Phi}_{c}=[{\phi}_{0}\,{\phi}_{1}\,{\phi}_{2}\,{\phi}_{3}\,{\phi}_{4}]$ illustrated by (a) five repeats of the primitive cell periodic state. (b) An internal perturbation $h_{z}$ has the same periodicity as the periodic state. Its spectrum, evaluated in Sec. VIII, is discrete. (c) A transverse perturbation $h_{z}$ is an arbitrary, aperiodic function over the infinite lattice. Its spectrum, evaluated by the Floquet-Bloch theorem in Sec. IX, is a continuous function of wave number $k$. Horizontal: lattice sites labelled by $z\in\mathbb{Z}$. Vertical: (a) value of field $\phi_{z}$, (b-c) perturbation $h_{z}$, plotted as a bar centred at lattice site $z$. Values of the field and perturbation are shown in blue within the primitive cell, and in orange outside the primitive cell.
@@
Sec. VIII Primitive cell stability.
@@
Sec. IX Bravais lattice stability.
A crystallographer or a field theorist starts –ab initio– with an infinite lattice or continuous spacetime, as in Fig. 8 (c). This, we claim in the introduction, Eq. (2), is the correct approach which –as we show here, Eq. (111)– yields (multi)periodic state weights that are multiplicative for repeats of spatiotemporally periodic solutions. The stability exponent per unit spacetime volume is the spacetime generalization of the temporal periodic orbit Lyapunov exponent, the mean instability per unit time. No matter what repeat of a prime periodic state one starts with, its stability exponent is always given by the same integral over the prime orbit Brillouin zone. From this follows the main result of our paper, the spatiotemporal zeta function of Sec. X.
(a)    Refer to caption
(b)    Refer to caption
Figure 9: (Color online)    One-dimensional lattice orbit Jacobian operator spectra, as functions of the reciprocal lattice wave number $k$. For time-reversal invariant systems the spectra are $k\to-k$ symmetric. (a) The steady state $\Lambda(k)$ spectrum Eq. (90), as a function of the first Brillouin zone wave number $k\in(-\pi,\pi]$, plotted for $\mu^{2}=1$ value. Any period-$n$ primitive cell Eq. (65) orbit Jacobian matrix spectrum consists of $n$ discrete points embedded into $\Lambda(k)$, for example period-3 (red triangles) and period-4 (magenta diamonds) eigenvalues. (b) The nonlinear $\phi^{3}$ theory $\Lambda_{01,\pm}(k)$ spectrum Eq. (115) of the Bravais lattice $\mathcal{L}_{01}$ tiled by the period-2 periodic state $\mathsf{\Phi}_{01}=\{\phi_{0},\phi_{1}\}$, together with the eigenvalues of 3rd repeat (red triangles) and 4th repeat (magenta diamonds) primitive cells. Plotted for $\mu^{2}=5$ value. See Appendix C.1. From companion paper I. [16]
VIII Primitive cell stability
As we now explain, it is crucial that we distinguish the finite primitive cell orbit Jacobian matrix (finite volume primitive cell stability, discussed in this section) from the infinite orbit Jacobian operator (infinite Bravais lattice stability, discussed in Sec. IX) in stability calculations.
To the best of our knowledge, in all current implementations of the periodic orbit theory, [18, 19, 45, 136, 20] the calculations are always carried out on finite primitive cells, so a ‘chaos’ expert is free to skim over this section - it is a recapitulation of Hénon, Lorentz, etc., calculations in the spatiotemporal, field theoretic language. The radical departure takes place in Sec. IX.
@@ VIII Primitive cell stability
VIII.1 Primitive cell steady state stability in one dimension
@@
Example: The spectrum of orbit Jacobian matrix for a steady state of period-3.
@@
VIII.2 Primitive cell steady state stability in two dimensions
@@
$p_{j}$ is the lattice momentum in $j$th direction, and $L_{j}$} is the period of the primitive cell $\mathbb{A}$ in $j$th direction, with ${N}_{\mathbb{A}}$ orbit Jacobian matrix eigenvalues $\Lambda_{m}=\mathsf{p}_{m}^{2}+\mu^{2}$ taking values on the reciprocal lattice sites $\mathsf{k}$, indexed by integer multiplets $m=m_{1}m_{2}\cdots{m_{d}}$. The inverse $1/(\mathsf{p}^{2}+\mu^{2})$ is known as the free-field propagator.
@@
Figure 10: (Color online) (a) As in Fig. 7 (b): Translations of reciprocal primitive vectors $\tilde{\mathbf{a}}_{1}$ and $\tilde{\mathbf{a}}_{2}$ (see Eq. (80), Eq. (89)) generate the reciprocal lattice $\mathcal{L}_{\tilde{\mathbb{A}}}$ (red points), with (shaded) the reciprocal primitive cell $\tilde{\mathbb{A}}$. (b) By convention, one restricts the range of wave numbers to (shaded) the first Brillouin zone $\mathbb{B}$, with $k_{1},k_{2}\in(-\pi,\pi]$.
Example: The spectrum of steady state orbit Jacobian matrix, $3\!\times\!2]_{1}$ primitive cell.
Consider primitive cell $3\!\times\!2]_{1}$} of example Eq. (88), drawn in Fig. 7 (a). The screw boundary condition yields S/T=1/2{S}/{T}=1/2. The wave numbers $\mathsf{k}$ in Eq. (94) are indexed by integer pairs $m_{1}=0,1,2$ and $m_{2}=0,1$. The $\mathsf{p}_{m_{1}m_{2}}^{2}$ in the reciprocal lattice orbit Jacobian matrix Eq. (96) is
@@
with, for example, the $(\tilde{{\mathcal{J}}}_{\mathbb{A}})_{21}$ eigenvalue $\Lambda_{21}=4+\mu^{2}$, and so on. Figure 11 (a) offers a perspective visualization of stability eigenvalues over such reciprocal cell, in that case $\mathcal{L}_{[8\!\times\!8]_{0}}$ periodic state. The corresponding orbit Jacobians are products of the ${\mathcal{J}}_{\mathbb{A}}$ eigenvalues, some of which are tabulated in Tables 1 and 2.
@@
The values of the lattice momentum square happen to be integers only for the few smallest primitive cells: in general their values are expressed in terms of Hipparchus cord functions $\mathrm{crd}(2\pi\,m_{j}/{L_{j}})$, or what we here call ‘lattice momenta’ Eq. (91). However, for integer values of spatiotemporal cat Klein-Gordon mass-square $\mu^{2}$, the orbit Jacobians take integer values, so if we are not interested in details of the spectrum, their direct evaluation might be preferable. That we do in Appendix B.3.
@@
VIII.3 Primitive cell periodic state stability
@@
The one spatiotemporal dimension $\phi^{3}$ theory Eq. (45) has one period-2 prime orbit $\{\mathsf{\Phi}_{01},\mathsf{\Phi}_{10}\}$ (for details, see Appendix C.1) with 2 orbit Jacobian matrix Eq. (57) eigenvalues
@@
This confirms the assertion we had made in the introduction, Eq. (1): orbit Jacobians of primitive cell periodic states are not multiplicative for their repeats. (Continued in Sec. IX.2.)
@@
In Appendix C.2 we work out as a further example the primitive cell stability of two-dimensional $\phi^{4}$ theory Eq. (46) $2\!\times\!1]_{0}$} periodic state. The eigenvalues of the primitive cell prime periodic state and its repetition are plotted in Fig. 11 (b). Next, note that the primitive cell of Bravais lattice $6\!\times\!4]_{0}$ can be tiled by twelve repeats of a prime $2\!\times\!1]_{0}$ periodic state. The eigenvalues of its orbit Jacobian matrix, plotted in Fig. 11 (b), lie on the two orbit Jacobian operator Bloch bands, located at twelve wave vectors in the first Brillouin zone of $6\!\times\!4]_{0}$: $k_{1}=-\pi/3,0,\pi/3$ and $k_{2}=-\pi/2,0,\pi/2,\pi$. (Continued in Sec. IX.2.)
@@
IX Bravais lattice stability
In Sec. IV.1 we have defined the stability exponent of a periodic state over a finite volume primitive cell, and in Sec. VIII we have explained how to compute them, setting the stage for the main result of section, the reciprocal lattice evaluation of the stability exponent for the orbit Jacobian operator.
@@ IX Bravais lattice stability
Continuing on the discussion of Sec. VIII.3: $k_{j}=0$ eigenvalues correspond to ‘internal’ eigenstates, states of the same periodicity as the periodic state $\mathsf{\Phi}_{\mathbb{A}}$, evaluated here in Sec. VIII for the primitive cell $\mathbb{A}$. The $k_{j}\neq 0$ continuum corresponds to ‘transverse’ eigenstates, perturbations that exit the $\mathbb{A}$-symmetric subspace, as in primitive cell example Eq. (102).
In textbook arguments leading to the Bloch theorem, one notes that larger and larger spatiotemporal primitive cells correspond to denser and denser reciprocal primitive cells (see, for example, Fig. 11 (b), and the arguments of the next section), leading in the infinite primitive cell limit to a parametrization by continuum values of wave vectors. Here we always evaluate stability exponents on infinite Bravais lattices, as integrals over the first Brillouin zone.
IX.1 Steady state stability
@@
The steady state calculations are so simple, as their orbit Jacobians are fully diagonalized by discrete Fourier transforms. For a steady state the unit hypercube primitive cell state is prime, all other periodic states over larger primitive cells are non-prime repeats of the unit hypercube periodic state (see Sec. I.3).
IX.2 Periodic state stability
As discussed in Sec. VIII.3, the nonlinear field theory orbit Jacobian operators typically depend on the periodic state and cannot be diagonalized by discrete Fourier transforms. The eigenvalues of the orbit Jacobian matrices in the primitive cells of prime periodic states can only be computed numerically.

### v3

@@ VIII Reciprocal lattice
Assembling the reciprocal primitive vectors $\{\tilde{\mathbf{a}}_{j}\}$ into columns of the $d\!\times\!d]$ reciprocal primitive cell matrix $\tilde{\mathbb{A}}=\left[\tilde{\mathbf{a}}_{1},\tilde{\mathbf{a}}_{2},\cdots,\tilde{\mathbf{a}}_{d}\right],$ the reciprocity condition Eq. (80) takes the form
@@
VIII.1 Reciprocal primitive cell in one and two dimensions
@@
with eigenvalues $\{{e}^{\mathrm{i}{k}}\}$ being the n{n}-th roots of unity, indexed by integers $m$,
@@
Figure 9: (a) The intersection points $z$ of the light gray lines form the integer square lattice Eq. (6). The primitive vectors $\mathbf{a}_{1}=(3,0)$ and $\mathbf{a}_{2}=(1,2)$ form the primitive cell $\mathbb{A}={[3\!\times\!2]_{1}}$ (see Eq. (76) and Fig. 2 (a)), whose translations tile the Bravais lattice $\mathcal{L}_{\mathbb{A}}$ (red points). (b) The intersection points $k$ of the light gray lines form the reciprocal square lattice. Translations of reciprocal primitive vectors $\tilde{\mathbf{a}}_{1}$ and $\tilde{\mathbf{a}}_{2}$ (see Eq. (80), Eq. (89)) generate the reciprocal lattice $\mathcal{L}_{\tilde{\mathbb{A}}}$ (red points). (Shaded) The reciprocal primitive cell $\tilde{\mathbb{A}}$. A wave vector outside this region is equivalent to a wave vector within it by a reciprocal lattice translation. Note that the number of lattice sites within the reciprocal primitive cell $\tilde{\mathbb{A}}$ equals the number of sites within the spatiotemporal primitive cell $\mathbb{A}$.
@@
Sec. IX Primitive cell stability.
@@
Sec. X Bravais lattice stability.
A crystallographer or a field theorist starts –ab initio– with an infinite lattice or continuous spacetime, as in Fig. 1 (c). This, we claim in the introduction, Eq. (2), is the correct approach which –as we show here, Eq. (111)– yields (multi)periodic state weights that are multiplicative for repeats of spatiotemporally periodic solutions. The stability exponent per unit spacetime volume is the spacetime generalization of the temporal periodic orbit Lyapunov exponent, the mean instability per unit time. No matter what repeat of a prime periodic state one starts with, its stability exponent is always given by the same integral over the prime orbit Brillouin zone. From this follows the main result of our paper, the spatiotemporal zeta function of Sec. XI.
IX Primitive cell stability
As we now explain, it is crucial that we distinguish the finite primitive cell orbit Jacobian matrix (finite volume primitive cell stability, discussed in this section) from the infinite orbit Jacobian operator (infinite Bravais lattice stability, discussed in Sec. X) in stability calculations.
To the best of our knowledge, in all current implementations of the periodic orbit theory, [18, 19, 45, 142, 20] the calculations are always carried out on finite primitive cells, so a ‘chaos’ expert is free to skim over this section - it is a recapitulation of Hénon, Lorentz, etc., calculations in the spatiotemporal, field-theoretic language. The radical departure takes place in Sec. X.
@@ IX Primitive cell stability
IX.1 Primitive cell steady state stability in one dimension
@@
Example: The spectrum of an orbit Jacobian matrix for a steady state of period-3.
@@
IX.2 Primitive cell steady state stability in two dimensions
@@
where $p_{j}$ is the lattice momentum in the $j$th direction, and $L_{j}$} is the period of the primitive cell $\mathbb{A}$ in the $j$th direction, with ${N}_{\mathbb{A}}$ orbit Jacobian matrix eigenvalues $\Lambda_{m}=\mathsf{p}_{m}^{2}+\mu^{2}$ taking values on the reciprocal lattice sites $\mathsf{k}$, indexed by integer multiplets $m=m_{1}m_{2}\cdots{m_{d}}$. The inverse $1/(\mathsf{p}^{2}+\mu^{2})$ is known as the free-field propagator.
@@
Example: The spectrum of the steady state orbit Jacobian matrix, $3\!\times\!2]_{1}$ primitive cell.
Consider the primitive cell $3\!\times\!2]_{1}$} of example Eq. (88), drawn in Fig. 9 (a). The screw boundary condition yields S/T=1/2{S}/{T}=1/2. The wave numbers $\mathsf{k}$ in Eq. (94) are indexed by integer pairs $m_{1}=0,1,2$ and $m_{2}=0,1$. The $\mathsf{p}_{m_{1}m_{2}}^{2}$ in the reciprocal lattice orbit Jacobian matrix Eq. (96) is
@@
with, for example, being the $(\tilde{{\mathcal{J}}}_{\mathbb{A}})_{21}$ eigenvalue $\Lambda_{21}=4+\mu^{2}$, and so on. Figure 10 (a) offers a perspective visualization of stability eigenvalues over such reciprocal cell, in that case $\mathcal{L}_{[8\!\times\!8]_{0}}$ periodic state. The corresponding orbit Jacobians are products of the ${\mathcal{J}}_{\mathbb{A}}$ eigenvalues, some of which are tabulated in Tables 1 and 2.
(a)   Refer to caption
(b)   Refer to caption
Figure 10: Square spatiotemporal lattice orbit Jacobian operator spectra, as functions of the wave vectors $(k_{1},k_{2})$. For time and space-reflection and interchange invariant periodic states the spectra are $k_{1}\to-k_{1}$, $k_{2}\to-k_{2}$ and $k_{1}\leftrightarrow k_{2}$ symmetric. (a) The steady state $\Lambda(\mathsf{k})$ Bloch band Eq. (106) as a function of the wave vector $\mathsf{k}$, plotted for $\mu^{2}=1$ value. Black dots are eigenvalues of the orbit Jacobian matrix of periodic states over primitive cell with periodicity $8\!\times\!8]_{0}$. (b) The two-dimensional $\phi^{4}$ lattice field theory spectra of the Bravais lattice $\mathcal{L}_{[2\!\times\!1]_{0}}$ periodic state Eq. (198), plotted for $\mu^{2}=5$ value. Black dots are eigenvalues of the orbit Jacobian matrix of a $6\!\times\!4]_{0}$ primitive cell tiled by 12 repeats of a prime $2\!\times\!1]_{0}$ periodic state, with $\Lambda_{\pm}(\mathsf{k})$ Bloch bands computed in Appendix C.2.
(a)    Refer to caption
(b)    Refer to caption
Figure 11: One-dimensional lattice orbit Jacobian operator spectra, as functions of the reciprocal lattice wave number $k$. For time-reversal invariant systems the spectra are $k\to-k$ symmetric. (a) The steady state $\Lambda(k)$ spectrum Eq. (90), as a function of the first Brillouin zone wave number $k\in(-\pi,\pi]$, plotted for the $\mu^{2}=1$ value. Any period-$n$ primitive cell Eq. (65) orbit Jacobian matrix spectrum consists of $n$ discrete points embedded into $\Lambda(k)$, for example period-3 (red triangles) and period-4 (magenta diamonds) eigenvalues. (b) The nonlinear $\phi^{3}$ theory $\Lambda_{01,\pm}(k)$ spectrum Eq. (115) of the Bravais lattice $\mathcal{L}_{01}$ tiled by the period-2 periodic state $\mathsf{\Phi}_{01}=\{\phi_{0},\phi_{1}\}$, together with the eigenvalues of third repeat (red triangles) and fourth repeat (magenta diamonds) primitive cells. Plotted for the $\mu^{2}=5$ value. See Appendix C.1. From companion paper I. [16]
@@
The values of the lattice momentum square happen to be integers only for the few smallest primitive cells: in general their values are expressed in terms of Hipparchus cord functions $\mathrm{crd}(2\pi\,m_{j}/{L_{j}})$, or what we here call ‘lattice momenta’ Eq. (91). However, for integer values of spatiotemporal cat Klein-Gordon mass-square $\mu^{2}$, the orbit Jacobians take integer values, so if we are not interested in details of the spectrum, their direct evaluation might be preferable, and that we do in Appendix B.3.
@@
IX.3 Primitive cell periodic state stability
@@
The one spatiotemporal dimension $\phi^{3}$ theory Eq. (45) has one period-2 prime orbit $\{\mathsf{\Phi}_{01},\mathsf{\Phi}_{10}\}$ (for details, see Appendix C.1) with two orbit Jacobian matrix Eq. (57) eigenvalues
@@
This confirms the assertion we had made in Sec. I, Eq. (1): orbit Jacobians of primitive cell periodic states are not multiplicative for their repeats. This will be discussed further in Sec. X.2.
@@
In Appendix C.2 we work out as a further example the primitive cell stability of two-dimensional $\phi^{4}$ theory Eq. (46) $2\!\times\!1]_{0}$} periodic state. The eigenvalues of the primitive cell prime periodic state and its repetition are plotted in Fig. 10 (b). Next, note that the primitive cell of Bravais lattice $6\!\times\!4]_{0}$ can be tiled by twelve repeats of a prime $2\!\times\!1]_{0}$ periodic state. The eigenvalues of its orbit Jacobian matrix, plotted in Fig. 10 (b), lie on the two orbit Jacobian operator Bloch bands, located at twelve wave vectors in the first Brillouin zone of $6\!\times\!4]_{0}$: $k_{1}=-\pi/3,0,\pi/3$ and $k_{2}=-\pi/2,0,\pi/2,\pi$. This will be discussed further in Sec. X.2.
@@
X Bravais stability
In Sec. V.1 we have defined the stability exponent of a periodic state over a finite volume primitive cell, and in Sec. IX we have explained how to compute them, setting the stage for the main result of section, the reciprocal lattice evaluation of the stability exponent for the orbit Jacobian operator.
@@ X Bravais stability
Continuing on the discussion of Sec. IX.3: $k_{j}=0$ eigenvalues correspond to ‘internal’ eigenstates, states of the same periodicity as the periodic state $\mathsf{\Phi}_{\mathbb{A}}$, evaluated here in Sec. IX for the primitive cell $\mathbb{A}$. The $k_{j}\neq 0$ continuum corresponds to ‘transverse’ eigenstates, perturbations that exit the $\mathbb{A}$-symmetric subspace, as in primitive cell example Eq. (102).
In textbook arguments leading to the Bloch theorem, one notes that larger and larger spatiotemporal primitive cells correspond to denser and denser reciprocal primitive cells (see, for example, Fig. 10 (b), and the arguments of the next section), leading in the infinite primitive cell limit to a parameterization by continuum values of wave vectors. Here we always evaluate stability exponents on infinite Bravais lattices, as integrals over the first Brillouin zone.
X.1 Steady state stability
@@
Figure 12: (a) As in Fig. 9 (b): Translations of reciprocal primitive vectors $\tilde{\mathbf{a}}_{1}$ and $\tilde{\mathbf{a}}_{2}$ (see Eq. (80), Eq. (89)) generate the reciprocal lattice $\mathcal{L}_{\tilde{\mathbb{A}}}$ (red points), with (shaded) the reciprocal primitive cell $\tilde{\mathbb{A}}$. (b) By convention, one restricts the range of wave numbers to (shaded) the first Brillouin zone $\mathbb{B}$, with $k_{1},k_{2}\in(-\pi,\pi]$.
@@
The steady state calculations are so simple, as their orbit Jacobians are fully diagonalized by discrete Fourier transforms. For a steady state the unit hypercube primitive cell state is prime, all other periodic states over larger primitive cells are non-prime repeats of the unit hypercube periodic state (see Sec. II.3).
X.2 Periodic state stability
As discussed in Sec. IX.3, the nonlinear field theory orbit Jacobian operators typically depend on the periodic state and cannot be diagonalized by discrete Fourier transforms. The eigenvalues of the orbit Jacobian matrices in the primitive cells of prime periodic states can only be computed numerically.